In a survey, 13 people were asked how much they spent on their child's last birthday gift. The results were roughly bell-shaped with a mean of $39 and standard deviation of $18. Estimate how much a typical parent would spend on their child's birthday gift (use a 95% confidence level).

Give your answers to one decimal place. Provide the point estimate and margin or error.
Answer #1

Answer:

Given,

n = 13

xbar = 39

s = 18

degree of freedom = n - 1

= 13 - 1

= 12

For 95% confidence interval , t(0.025 , 12) = 2.179

Interval = (xbar +/- t*s/sqrt(n))

= 39 +/- 2.179*18/sqrt(13)

= 39 +/-10.9

= (39 - 10.9 , 39 + 10.9)

= (28.1 , 49.9)

Point estimate = xbar = 39

Margin of error = t*s/sqrt(n)

= 2.179*18/sqrt(13)

= 10.9